Mission Operations Days: 2006/301 to 2006/307 October 28, 2006 00:00:00 GMT to November 03, 2006 00:00:00 GMT
Terra Spacecraft and MODIS Instrument Status:
|AM-1 "Terra" is in Normal Mode PFM MODIS is in A-Side Nominal Science Mode (Cross Strapped with FR B)|
|Blackbody||A On (290K); B Off||Nominal Science Mode|
|Calibration Electronics||A On; B Off||Nominal Science Mode|
|Control Processor||A On; B Off||Nominal Science Mode|
|Nadir Door||Unlatched, Open||Nominal Science Mode|
|Space View Door||Unlatched, Open||Nominal Science Mode|
|Solar Diffuser Door||Unlatched, Screened||Permanent Configuration|
|FDDI Formatter||A On; B Off||Nominal Science Mode|
|FDDI Port||A On; B Off||Nominal Science Mode|
|FIFO Memory||#1 & #2 On; #3 & #4 Off||Nominal Science Mode|
|Format Processor||A Off; B On||Nominal Science Mode|
|Power Supplies||#1 On, #2 Off||Nominal Science Mode|
|PV VIS FPA||A On; B Off||Nominal Science Mode|
|PV NIR FPA||A On; B Off||Nominal Science Mode|
|PV SMIR FPA||A On; B Off||Nominal Science Mode|
|PV LWIR FPA||A On; B Off||Nominal Science Mode|
|PC LWIR FPA||A On; B Off||Nominal Science Mode|
|RC Outgas Heaters||All 3 (CS, IS, OS) Off||Nominal Science Mode|
|RC LWIR FPA Heater||On (83K)||Nominal Science Mode|
|RC SMIR FPA Heater||Off||Nominal Science Mode|
|Scan Assembly||A On; B Off||Nominal Science Mode|
|SDSM||Off||Nominal Science Mode|
||Nominal Science Mode|
|PS1 Survival Heater||Enabled||Nominal Science Mode|
|PS2 Survival Heater||Enabled||Nominal Science Mode|
|Timing Generator||A On; B Off||Nominal Science Mode|
|Flight Software||Rev BD + 11 sets of patches
||Nominal Science Mode|
|Inhibit Ids Set||none||Nominal Science Mode|
|TMON 66||Disabled||Nominal Science Mode|
|TMON 67||Disabled||Nominal Science Mode|
|TMON 68||Enabled||Nominal Science Mode|
|TMON 69||Disabled||Nominal Science Mode|
This Week's Completed MODIS Activities:
Saturday, October 28, 2006
Sunday, October 29, 2006
Monday, October 30, 2006
Tuesday, October 31, 2006
2006/304 21:06:09-21:14:57 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement.
Wednesday, November 1, 2006
Thursday, November 2, 2006
Friday, November 3, 2006
This Week's Scheduled MODIS Activities Not Completed:
Upcoming MODIS Events:
Calibrations for November will include a biweekly SDSM, a monthly SRCA Full Radiometric, and a quarterly SRCA Full Spatial.
Thursday, November 9, 2006 Lunar Roll Maneuver (-13.28 deg roll) 2006/313 16:58:14 ATC MOD_OA03_FR_NIGHT_RATE (03:23 early) 2006/313 17:35:31-17:42:17 ATC MOD_OA08_ROLL_MAN 2006/313 17:54:16 ATC MOD_OA03_FR_DAY_RATE (03:23 late)
The next Terra MODIS Lunar Calibration Roll will be on DOY 2006/313 (November 9, 2006). [Previous reports stated 2006/315, which was incorrect.] The roll angle will be -13.2814 degrees. Data collect will be from 17:35:39-17:42:09.
The next two Terra Inclination Adjustment Maneuvers (IAMs) will occur on November 14 and 21 - DOY 2006/318 and 2006/325, respectively. [NOTE: the date of the second maneuver changed since the last weekly report.] If needed, a back-up IAM would be performed on 2006/334. The times of the maneuvers and the MODIS contingency activities are tentatively scheduled as follows:
Tuesday, November 14, 2006 2006/318 15:46-16:21 Terra Inclination Adjustment Maneuver #16 (planned - not yet scheduled) 2006/318 16:21:38-16:23:19 ATC MOD_CLOSE_NAD_SVD.1 IAM Contingency: Closes the NAD and SVD 100 seconds prior to S/C Day Note: This activity will not execute if the IAM is successful.
Tuesday, November 21, 2006 2006/325 15:52-16:28 Terra Inclination Adjustment Maneuver #17 (planned - not yet scheduled) 2006/325 16:28:19-16:30:00 ATC MOD_CLOSE_NAD_SVD.1 IAM Contingency: Closes the NAD and SVD 100 seconds prior to S/C Day Note: This activity will not execute if the IAM is successful.
General Instrument Comments:
MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.
MODIS Telemetry Trends:
Nominal telemetry trending is ongoing.
Non-MODIS Significant Events:
The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.
Limited Life Item Status:
SRCA 10W Lamp #1: 256.9 of 500 hours SRCA 10W Lamp #2: 172.1 of 500 hours SRCA 10W Lamp #3: 190.3 of 500 hours SRCA 10W Lamp #4: 81.8 of 500 hours
SRCA 1W Lamp #1: 570.5 of 4000 hours SRCA 1W Lamp #2: 278.7 of 4000 hours
Solar Diffuser Door: 2146 of 3022 Movements Nadir Aperture Door: 540 of 1316 Movements Space View Door: 443 of 1316 Movements